Suppose you have $600 in your account. You bought 2 shares of Apple stock (AAPL) at $204/share and 1 share of Microsoft stock (MSFT) at $102/share.

The weight of AAPL in your portfolio is:

The weight of MSFT in your portfolio is:

The weight of Cash in your portfolio is:
Answer #1

Amount in AAPL=2*204=408
Amount in MSFT=102
Cash Left=600-408-102=90

Weight of AAPL=408/600=68%
Weight of MSFT=102/600=17%
Weight of CASH=90/600=15%